One number is 3 less than another. If 10 times the smaller number minus 4 times the larger number is 24, find the two numbers.
Found 2 solutions by mathman409, checkley75:
Answer by mathman409(67)

6 and 9 are your answers. If you want to write an algebraic equation, you can, but for small numbers like these it is faster to use guess and check.

Answer by checkley75(3666)

LET X=THE SMALLER NUMBER & X+3 THE LARGER NUMBER THEN:
10X-4(X+3)=24
10X-4X-12=24
6X=24+12
6X=36
X=36/6
X=6 ANSWER FOR THE SMALLER NUMBER THEN:
6+3=9 FOR THE LARGER NUMBER.
PROOF
10*6-4(9)=24
60-36=24
24=24